# Inferring Capabilities from Task Performance with Bayesian Triangulation

### v1

@@ Abstract
As machine learning models become more general, we need to characterise them in richer, more meaningful ways. We describe a method to infer the cognitive profile of a system from diverse experimental data. To do so, we introduce measurement layouts that model how task-instance features interact with system capabilities to affect performance. These features must be triangulated in complex ways to be able to infer capabilities from non-populational data – a challenge for traditional psychometric and inferential tools. Using the Bayesian probabilistic programming library PyMC, we infer different cognitive profiles for agents in two scenarios: 68 actual contestants in the AnimalAI Olympics and 30 synthetic agents for O-PIAAGETS, an object permanence battery. We showcase the potential for capability-oriented evaluation.
@@ 1 Introduction
What does success or failure tell us about a system’s capabilities? In isolation, performance results are hard to interpret: most interesting tasks require several different abilities to be successfully combined. instances. For example, an agent exploring a 3D environment to find a reward needs to understand that the reward still exists when occluded (object permanence). It must then remember where it is, and successfully navigate to it. If the agent fails, we cannot know the cause: lack of object permanence, limited memory or navigation skills. This leads to poor predictability about the performance of the agent for new tasks, in or out of the training distribution.
Figure 1: This task instance requires the agent to: (a) understand that the reward still exists when occluded (object permanence), (b) remember which of the three walls it is occluded behind while it navigates towards it (memory), and (c) reach it over ramps and around the wall (navigation).
Figure 2: Triangulation in the measurement layouts using results from task instances like this one. Bottom-up inference from three tasks (in green and red for success and failure respectively) leading to the cognitive profile. Top-down inference (in blue) predicting failure for the fourth task.
@@
However, suppose we also knew that the agent performs well on tasks that involve complex navigation, and the same for tasks involving memory. We could now infer that the system’s failure is likely a failure of object permanence by triangulation, a robust inferential method in formal epistemology (Heesen, Bright, and Zucker 2019) and causal (Bayesian) reasoning.
@@
This bottom-up inference is illustrated in green and red in Fig. 2. We can use this information to draw conclusions about where the system is safe to deploy, and to direct efforts to improve it. We can also predict behaviour in new scenarios: for example, the agent will likely fail at a new task with high object permanence demands, as illustrated in blue in Fig. 2.
@@
This example demonstrates that we can leverage the fact that different tasks—and their constituent instances—have different demands to help us extract conclusions about system capabilities from patterns of task performance. However, these conclusions require an analytical approach that facilitates complex inferences. If we were to examine each task or instance in isolation or to simply aggregate performance across all tasks, this triangulation of abilities would not be possible. To address this, we propose a novel Bayesian approach to evaluation that enables simultaneous inferences about multiple abilities from patterns of task performance. This approach has a series of principles facilitating a capability-oriented approach to the evaluation of machine learning models:
1. We do not simply break out performance by ‘categories’. Related work employs a ‘taxonomy of tasks’ (see e.g., Table 1 in (Liang et al. 2022)), each placed into one or more categories. This limits the inferential capacity and predictability of these category aggregates for new tasks. Instead, we evaluate based on instance demands.
2. Unlike psychometric approaches such as factor analysis (FA), structural equation modelling (SEM) or item response theory (IRT), we do not rely on populational data (Gustafsson and Undheim 1996). Instead, we infer the cognitive profile of a single subject from their performance data alone.
3. We do not ‘reify’ extracted latent variables as capabilities, which has been done hierarchically elsewhere, (e.g., the Cattell-Horn-Carroll model (Keith and Reynolds 2010)).
Instead, we identify general linking functions that help derive the requisite capabilities from the demands of the tasks.
@@
4. Probabilistic models currently used in cognitive modelling (Lee 2011) are often simple at the level of the individual and cannot convey the required probabilistic expressions connecting individual performance, capabilities and task features. Instead, complex interactions allow for Bayesian triangulation which we use for informative evaluation. In this paper we find an operational solution to these inferential challenges in the form of Measurement Layouts: specialised and semantically-rich hierarchical Bayesian networks. Using the inference power of the No U-Turn Sampler (Hoffman and Gelman 2014) in PyMC (Salvatier, Wiecki, and Fonnesbeck 2016), a probabilistic programming engine, we can triangulate from performance and the task’s cognitive demands to the capability and bias levels for each subject (their cognitive profile). Then, performance can be inferred downstream for new tasks.
We derive the cognitive profiles in two scenarios: the contestants in the AnimalAI Olympics (Crosby et al. 2020) and several synthetic agents for O-PIAAGETS (Voudouris et al. 2022), an object permanence battery. We show how systems with similar overall performance can have very different capability levels and biases.
2 Background
AI systems are often evaluated on one or more standard benchmarks. Because of the remarkably broad capabilities of modern systems, benchmarks are becoming larger and more diverse. For instance, BIG-bench (Srivastava et al. 2022) contains over 200 tasks for evaluating language models. Aggregating many tasks into a single score does not allow for nuanced conclusions about where or why a system fails (Burnell et al. 2023), and has little value for predicting system performance a specific unseen task.
Analysing 200 tasks separately would lead to an unwieldy and hard to interpret set of results. One option is the categorisation of tasks across several dimensions, providing a moderate level of insight into where a model succeeds or fails (Osband et al. 2019; Ilievski et al. 2021; Crosby et al. 2020; Liang et al. 2022). However, conclusions are highly dependent on how instances are assigned to categories, and predictive value is limited inside and across categories.
An widely-used approach that does not require any taxonomy of capabilities is to extract and reify them from the data. Given experimental data for many tasks and subjects, we can use Factor Analysis (FA) to infer those factors that best explain the observed variance in the population. Some structure can be inferred from FA, as in the classical hierarchical models (Cattell-Horn-Carroll model (Keith and Reynolds 2010)), but today it is common to use more powerful techniques such as Structural Equation Modelling (SEM) (Ullman and Bentler 2012). However, SEM assumes linear relationships between variables and that dependent variables have normally distributed sample means. These assumptions are unlikely in the real world. Further, SEM is a populational technique requiring a large sample of individuals in order to confirm structural relationships.
A more granular approach is to identify the difficulty of each task instance and the ability of each system based on the relationship between task difficulty and the system’s performance. This idea is the cornerstone of Item Response Theory (IRT) (Embretson and Reise 2000; De Ayala 2009).
When several abilities exist, we must rely on multidimensional IRT(Reckase 2009) However, IRT is also populational, and derives abilities and difficulties relative to population averages, therefore the parameters of the same item can change if we add new items, or the abilities of a system may change if we add more systems. This poses a problem for AI evaluation, where the population of individual systems is constantly changing and advancing.
Still, with FA, SEM, IRT or other techniques, patterns of performance are mapped to capabilities, making up a multi-dimensional characterisation of the subject that is sometimes called a cognitive profile (Letteri 1980). Cognitive profiles make it possible to identify differences that might go unnoticed in aggregated scores, make predictions about future performance, and gain insights into the relationships between different abilities (e.g., deficits in mental rotation are one cause of poor spatial navigation).
Most psychological assessment is populational. Exceptions can mostly be found in developmental psychology and animal cognition (Moran et al. 2017; Bergman and Andersson 2015; Shaw and Schmelz 2017). The scarcity of subjects and the difficulty of experiments results in an emphasis on carefully designed tasks that ensure a systematic coverage of the relevant features and possible distractors. On the other hand, if the cognitive mechanisms for solving the task are accurate and enough variation is introduced, as in the example of Fig. 2, the results can be sufficient to estimate the capabilities of a single individual, independently of the other evaluated individuals. This is the type of approach that we need to evaluate ML models and AI systems.
Fortunately, there is recent movement towards producing training and test variations, augmented examples or corruptions to test systems, through adversarial learning (Zellers et al. 2018; Nie et al. 2019; Rozen et al. 2019; Kiela et al. 2021), procedural content generation (Risi and Togelius 2020) and other domain-specific techniques (Sugawara, Stenetorp, and Aizawa 2020). These variations have also been inspired by software testing, such as (Marijan and Gotlieb 2020) and (Ribeiro et al. 2020). Some of this effort results in annotated datasets, especially for reinforcement learning — from the inspiration behind the AnimalAI Olympics (Crosby et al. 2020) to the fully-systematic O-PIAAGETS (Voudouris et al. 2022). However, the next step of using these annotated data for capability-oriented evaluation has not been taken yet.
Not using populational data has to be compensated by a larger number of varied instances. When many dimensions of variation are introduced, the inference problem becomes complex. Traditional Bayesian Networks (BN) (Ben-Gal 2008) can be used for arbitrarily sophisticated probability distributions. For complex networks, though, performing the inference calculation is computationally intractable (Cooper 1990). Approximation techniques for Bayesian inference (BI) have been introduced, including the Markov Chain Monte Carlo family (Andrieu et al. 2003), allowing for the inference of latent values in more complex BNs, particularly those with multiple levels of dependency and complex relations between variables. Taking advantage of these recent advances in Bayesian approximation, evaluation procedures, and computing power allows us to extend the benefits of capability-oriented evaluation to machine learning for the first time, enabling more robust, thorough, and principled evaluations of ML systems.
@@ 3 Measurement Layouts
Performance can be understood as a function of task demands and capability levels. The precise relation between the two depends on the task features and the tested capabilities. For instance, in a memory problem, the number of objects to remember in a particular task demands a memory capability level at least as high as the number of objects. However, the difference between task demands and capability levels is likely not sufficient to account for all variance in observed performance. Cognitive biases, such as a preference or aversion to apparently irrelevant features like colour, can also affect performance. Finally, AI systems can exhibit a lack of robustness caused by some unaccounted factors or noise. Cognitive profiles need to include all of these. Let us define more precisely a task characterisation and the corresponding cognitive profile. We will then detail how they are connected by the measurement layouts.
Definition 3.1.
A task characterisation is a set of observable, usually constructed, meta-features $X$, expressing cognitive demands and other high-level properties of the task.
@@
For instance, in an image classification problem, the level of blur of the image or the level of clutter of the background could constitute perceptual demands. Many other properties may also, in practice, affect performance.
Definition 3.2.
The cognitive profile of a system is a tuple $\langle C,B,R\rangle$. Where $C$, $B$, and $R$ are vectors of capability levels, bias, and robustness values respectively.
@@
Capability levels represent what the system can do, bias values represent some other preferences or limitations that may impact performance in a less monotonic way, and robustness levels account for reliability issues, unexplained or random effects (noise) on either the agent or environment.
Definition 3.3.
A measurement layout is a directed acyclic graph that connects, through linking functions, the meta-features of a task characterisation with the cognitive profile of a system, in order to predict observed performance.
@@
We define measurement layouts as Hierarchical Bayesian Networks (HBN) (Murphy 2023) with the following characteristics: Within the measurement layout we include a node for every meta-feature in every task instance and a node for each element of the system’s cognitive profile. All of them are roots of the graph. A connection $A\rightarrow B$ denotes that $B$ is conditionally dependent on $A$. Linking functions are used to formally define relationships between dependent nodes.
All nodes within the HBN define a probability distribution. We can classify nodes into three categories depending on the provenance of their summary statistics:
Metafeatures: These come from the task characterisation and are fixed, observable values. They are represented as a fixed Dirac-delta distribution.
Cognitive Profile Nodes: These are elements of the cognitive profile relevant to assessed task’s demands. This includes required capabilities and likely biasses. These often combine with meta-features and feed in to derived nodes. The range and distribution of these elements derives from the task characteristics. To make our inferences as data-driven as possible these begin with uninformative priors.
Derived Nodes: These are instance-level inferences. They are combinations of meta-features and/or cognitive profile elements that have their parameters derived from dependent nodes. Information to compute these values is encoded in the linking function for the node, the output of which is used as the summary statistics of the node’s distribution.
One derived node is the observable performance for the task—the only leaf node—but most are Intermediate Non-Observable Nodes (INON), which represent intermediate performances or effects. All three node types were seen in Fig. 2 (right). The connective structure can be enriched by domain knowledge determining which cognitive profile elements and meta-features are expected to affect the values of instance-level inference nodes. These relationships can be higher-order and complex, with certain INON nodes requiring the nuanced confluence of many cognitive profile elements and meta-features. Complex connective structure (and the requirement for increased levels of domain knowledge) can arise from the need to accurately distinguish between success, failure, and the multitude of confounding behaviour that a carefully designed benchmark must account for.
By a linking function, we refer to any mathematical expression that maps values from the outputs of one node’s probability distribution to the input of another.
For instance, it could be a sigmoid function of the difference between capabilities and demands. It could also be a product of (non-compensatory) capabilities, representing a quantitative and differentiable expression for a logical ‘and’. Alternatively, it could be some additive linear equation, representing a quantitative expression of a logical ‘or’, when some preceding nodes are compensatory—or a generalised mean in between. The functions can scale the incoming nodes, so that we adjust the ranges for capabilities, biases and noise. The possibilities are endless; in practice we use only a few kinds of linking functions. We give the specific equations for the linking functions that we used used in this work in Tables 3 and 5 in the appendix. These are used in the two scenarios that follow.
@@
A key conceptual difference between our measurement layouts and typical use cases for HBNs is that our approach is trying to capture a hierarchical dependency relation on capabilities and demands, rather than encoding information on hyper-priors. The linking functions between nodes are often non-linear, conceived from domain knowledge of the task and applying ideas from cognitive modelling.
For instance, one of the key decisions is to determine when some capabilities are independent, and if so, whether they are compensatory or not, being realised with additive or multiplicative expressions respectively. If they are not independent, then an arrow has to connect them with a linking function expressing this dependency mathematically.
Given the cognitive profile of a system, and the characterisation of a new task, we can use the measurement layout to anticipate performance by forward (top-down) inference. But before that we need to infer the cognitive profile from the observed performance of other tasks by backward (bottom-up) Bayesian inference. For this, we utilise PyMC’s inference engine and the No U-Turn Sampler (Hoffman and Gelman 2014) approach to BI (see Appendix A for a detailed description). For this to accurately predict a system’s cognitive profile, the test battery
requires instances to jointly control for alternative explanations permitting performance results to ‘triangulate’ latent capabilities.
4 Cognitive Profiles for the AAI Olympics
To illustrate how a measurement layout identifies cognitive profiles, we first select a domain with simple task characterisations that can be straightforwardly related to cognitive profiles to predict performance.
The Animal-AI Olympics (AAIO) (Crosby et al. 2020) meets all these requirements. The AAIO was an open competition evaluating AI agents in a 3D environment, where agents had to get a reward by locating it and navigating in an arena.
Participants were asked to build systems (usually reinforcement learning agents) and train them on their own custom-built tasks in the AAI Environment. Screenshots for two tasks are shown in Fig. 7 located in Appendix B along with more information about AAIO and the environment. We select 69 tasks from the competition test set that evaluate simple goal-directed behaviour. We include 68 participant agents for which we know nothing about their design.
4.1 Measurement Layout
In this domain, following the analysis performed in (Burnell et al. 2022), the set of meta-features $X$ of the instances that characterise task instances are rewardSize, rewardDistance, Xpos and rewardBehind. These represent the reward’s size, distance, and if it is left, right, in front, or behind the agent.
We then compare the profile of the agent to the demands of the task to derive performance. This is our measurement layout, as can be seen in Fig. 4 (left). The metafeatures $X$ appear as observed (shaded) nodes in the model, appearing within the plate, with 69 instances. The elements of the cognitive profile appear as external latent (unshaded) variables.
The derived nodes lie at the intersection of the meta-features and the external cognitive profile nodes.
Figure 3: Measurement layout for the Animal AI Olympics.
Figure 4: Cognitive profiles of the 68 agents in AAIO, with names shown for some of them. The $x$-axis and $y$-axis show the inferred means for navigationAbility and visualAbility respectively, with their standard deviations as error bars in grey. The radius of each point represents the average performance, while the colour represents the noiseLevel, with red representing higher noise than blue.
Not visible in the figure are the linking functions associated with each arrow. The variable rightLeftEffect is calculated as the product between rightLeftBias and XPos, since left/right bias and XPos are represented with negative and positive values respectively, so the product gives a positive result when both go in the same direction. This rightLeftEffect, jointly with rewardBehind and rewardDistance affect navigationPerformance, a non-observable intermediate performance variable. The exact formulation of how these three variables are integrated can be found in Appendix C.2,
but broadly the linking function subtracts the effects of reward distance and position relative to the agent’s viewpoint (e.g., behind) from the navigation ability and adds the right-left bias to this. The linking function then applies a logistic function to this margin to give a value between 0 and 1 for navigationPerformance. The process to derive visualPerformance is similar, simply subtracting rewardSize ( inverted so the higher the smaller) from visualAbility.
Note that the ranges of these two abilities are given by the ranges of the task metafeatures, with navigationAbility in $[0,5.3]$ and with visualAbility in $[0,1.9]$ respectively, representing how distant and how big a reward—with their corresponding distance and size units—the agent can obtain. Finally, taskPerformance is Bernoulli-distributed on a product of navigationPerformance and visualPerformance adding the noiseLevel. We use the product because navigation and visual acuity are non-compensatory (i.e., navigation skills are of no use if the agent cannot see, and vice versa).
4.2 Inference and Analysis
Using the measurement layout and the 69 instances, we can use PyMC to infer a very simple cognitive profile for each of the 68 entrants of the competition. We run the NUTS inference algorithm with two sampling chains of size 1000 each, finding no inference divergences.
Fig. 4 (right) summarises the inferred capabilities and noise levels. We see the two main capabilities are correlated for this population, going from low performance in the bottom left area (e.g., ‘ice-play’ with 4% accuracy) to high performance in the top right area (e.g., ‘Trrrr’ with 99% accuracy). In these extremes, capabilities are nonexistent (close to 0) or saturated (to the maximum 5.3 and 1.9 respectively), as expected. But in the middle range, there are agents that have noticeable differences between their capabilities (e.g., ‘cocel’). Some have strong biases, such as ‘Thursday’ and ‘KoozytHiperdyne’ with left-right bias values of $1.14$ and $-0.78$ respectively (not shown on the plot). This can have an effect on navigation ability but not on visual ability.
The standard deviations of the abilities, shown as error bars, and the level of noise are useful to explain situations both when behaviour is conformant or not with the task characteristics. For instance, ‘Juohmaru’, with 54% accuracy, is a very conformant agent: performance decreases as rewards get smaller and further away, and when they are initially located behind from the agent. We see that navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, with a medium level of noise $0.26\pm 0.17$. But ‘sparklemotion’, with 36% accuracy, is very non-conformant, with abilities $0.40\pm 0.38$ and $0.59\pm 0.45$, and higher noise $0.31\pm 0.13$. Appendix C contains full characteristic grids of these two agents in Fig. 8 (demonstrating the conformant / non-conformant behaviour) and all inferred results in Fig. 9.
Note that we have inferred two different capabilities for each agent as well reliability measure of these estimates: how well they can predict performance (depending on the quality of the estimates, biases and remaining noise) for each particular instance given its meta-features. We do this without having different sub-batteries for each capability.
Our model can also predict performance on unseen tasks instances. To evaluate this, we randomly held out 20% of the task instances from the backwards inference stage and evaluated the predicted probabilities of success on them. Due to the small size of the dataset, we repeated this process 15 times and took the mean Brier score for each agent. Fig. 5(a) summarises these scores, which are lower than predictions based on aggregate performance for agents with success rates in the range $[0.2,0.7]$, exactly where there is variability to explain. Full results and discussion in Appendix E.
5 Cognitive Profiles for Object Permanence
Our analysis of the AAI Olympics used a relatively simple scenario as a proof of concept. But one limitation of this analysis is that we do not know the ”ground truth” of the agents’ abilities, so it is difficult to determine precisely how effective our approach was. It also remains unclear how well our approach works in complex scenarios with many abilities and metafeatures. Therefore, we next sought to demonstrate the scalability of our approach by analysing data from a more complex scenario—a rich test battery for evaluating Object Permanence. To ensure that we knew the ground truth abilities of the agents, we created a synthetic dataset by simulating the performance of 30 hypothetical agents based on specified cognitive profiles 111because of the complexity of the abilities that affect performance on this testbed, it would be extremely difficult to create a varied population of real agents with specific levels of abilities. This scenario tests whether our methods can recover complex cognitive profiles from performance data of agents for which we know the ground truth.
To do so, the authors of this paper were divided into two teams, Team A and Team B. Team A generated the synthetic dataset, while Team B defined the measurement layout and attempted to recover the cognitive profiles of the synthetic agents. Both teams were familiar with O-PIAAGETS and the general abilities, bias and elements that are combined in the battery. This two team approach allowed us to generate synthetic agents while keeping Team B blind to their profiles, removing the potential for bias and making the modelling effort more akin to a real-world scenario.
5.1 The Testbed and Synthetic Agents
We applied our measurement methodology to the Object-Permanence In Animal-AI: Generalisable Test Suites (O-PIAAGETS, (Voudouris et al. 2022)). Object Permanence (OP) is the understanding that objects continue to exist when they go out of view. Behaviourally, an agent has object permanence if they behave as though objects continue to exist even when they cannot directly perceive them. It is a key component of physical common-sense (Shanahan et al. 2020; Lake et al. 2017; Baillargeon et al. 2011), allowing biological agents to more successfully predict and interact with their environment. Object permanence is well studied in cognitive science, where several experimental designs are used to detect its presence in human and non-human animals (Scholl 2007; Flombaum and Scholl 2006; Hare and Tomasello 2005; Chiandetti and Vallortigara 2011). Many of these paradigms are represented in O-PIAAGETS, which contains over 13,000 instances. We selected three paradigms that test for allocentric object permanence, where objects are occluded independently of the observer’s actions. We also include basic control tasks within the test-suite to evaluate the agent’s non-OP capabilities and make the inference process more robust. Details about the test-bed design criteria, the included tasks, and why object permanence is difficult to identify are given in the appendix in D.1 and D.2.

### v2

@@ Abstract
As machine learning models become more general, we need to characterise their capabilities in richer, more interpretable ways that move beyond aggregated statistics on static benchmarks. We describe a method to infer the cognitive profile of a system from diverse experimental data. To do so, we introduce measurement layouts which model how task-instance features interact with system capabilities to explain performance. System capabilities can be estimated using Bayesian triangulation, inferring their value based on the performance of a model on tasks with different features. Our approach accurately recovers the cognitive profiles of hand-crafted behavioural agents, as well as estimating the cognitive profiles for deep reinforcement learning agents and human children in a virtual game environment. These cognitive profiles are significantly richer than aggregated benchmark statistics, summarising multiple distinct capabilities that explain behaviour, and are also significantly more predictive, accurately estimating performance on new, held-out tasks.
@@ 1 Introduction
What does success or failure tell us about a system’s capabilities? In isolation, performance results are hard to interpret: most interesting tasks require several different capabilities to be successfully combined. For example, consider an agent exploring a 3D environment similar to that in Figure 1(a), where in order for the agent to find the reward, it needs to understand that the reward still exists when occluded (known as object permanence). It must then remember where it is, and successfully navigate to it. If the agent fails on an instance of this task, the cause is unclear—is it a failure of object permanence, limited memory, or navigation skills? Without an answer to this question, it is difficult to predict performance for new tasks with different features. One way to approach this issue is to collect the performance of the agent on a benchmark of arbitrarily varied tasks (with and without occlusion, shorter or longer distances to the goal, etc.), and compute the proportion of successes across that set. This aggregated statistic approximates the expected probability of success on a new task instance, but it does so blind to its specific features. Moreover, this approach fails to explain why the agent passes or fails on a certain task.
(a) A three-cup instance. The agent must (a) retain the object’s existence when occluded, (b) remember which wall hides it, and (c) navigate around obstacles.
(b) Bayesian triangulation: bottom-up inference (green/red) yields a cognitive profile; top-down inference (blue) predicts performance on a new task.
Figure 1: (1(a)) shows a representative task instance; (1(b)) shows how Measurement Layouts infer and predict capabilities from instance-level results.
@@
Suppose, however, we also knew that the agent performs well on tasks that involve complex navigation, and the same for tasks involving memory. We could now infer, by triangulation, that the system’s failure is likely a failure of object permanence, a robust inferential method in formal epistemology (Heesen et al., 2019) and causal (Bayesian) reasoning.
@@
This bottom-up inference is illustrated in green and red in Fig. 1(b). We can use this information to draw conclusions about where the system is safe to deploy, and to direct efforts to improve it. We can also predict behaviour in new scenarios: for example, the agent will likely fail at a new task with high object permanence demands, as illustrated in blue in Fig. 1(b).
@@
This example demonstrates that we can leverage the fact that different tasks—and their constituent instances—have different demands to help us extract conclusions about system capabilities from patterns of task performance. However, these conclusions require an analytical approach that facilitates complex inferences. If we were to examine each task or instance in isolation or to simply aggregate performance across all tasks, this triangulation of abilities would not be possible. To address this, we propose a novel Bayesian approach to evaluation that enables simultaneous inferences about multiple capabilities from patterns of task performance. This approach is underpinned by several psychometric principles that we elucidate in Appendix A.
@@
In this paper we find an operational solution to these inferential challenges in the form of Measurement Layouts: specialised and semantically-rich hierarchical Bayesian networks. Using the inference power of the No U-Turn Sampler (Hoffman and Gelman, 2014) in PyMC (Salvatier et al., 2016a), a probabilistic programming engine, we can triangulate from performance and the task’s demands to the capability and bias levels for each subject (their cognitive profile). Then, performance can be inferred downstream for new tasks. Not only are cognitive profiles much more explanatory of what causes failure or success, but we find that they are also much more predictive than the traditional approach of using aggregated benchmark statistics to predict performance.
Our approach is agnostic to the task or type of system being evaluated, so long as there are instance-level task demands that can be related to distinct capabilities. To showcase our method, we construct measurement layouts and cognitive profiles for agentic systems in two scenarios: a simple navigation task and a more complex object permanence task. All code and data associated with our method can be found on GitHub here
@@ 2 Measurement Layouts
Performance can be understood as a function of task demands and capability levels. The precise relation between the two depends on the task features and the tested capabilities. For instance, in a memory problem, the number of objects to remember in a particular task demands a memory capability level at least as high as the number of objects. However, the difference between task demands and capability levels is likely not sufficient to account for all variance in observed performance. Cognitive biases, such as a preference or aversion to apparently irrelevant features like colour, can also affect performance. Finally, AI systems can exhibit a lack of robustness caused by some unaccounted factors or noise. Cognitive profiles need to include all of these.
@@
We define a task characterisation as the set of observable, usually constructed, meta-features $X$, expressing cognitive demands and other high-level properties of the task. For example, in an image classification problem, the level of blur of the image or the level of clutter of the background could constitute perceptual demands. Many other properties may also, in practice, affect performance.
@@
Similarly, we characterise the cognitive profile of an AI system as a tuple $\langle C,B,R\rangle$. Where $C$, $B$, and $R$ are vectors of capability levels, bias, and robustness values. Capability levels represent what the system can do, bias values represent some other preferences or limitations that may impact performance in a less monotonic way, and robustness levels account for reliability issues, unexplained or random effects (noise).
@@
We define a measurement layout as a directed acyclic graph that connects, through linking functions, the meta-features of a task characterisation with the cognitive profile of a system, in order to predict observed performance. We implement measurement layouts as Hierarchical Bayesian Networks (HBN) (Murphy, 2023) with the following characteristics: Within the measurement layout we include a node for every meta-feature in every task instance and a node for each element of the system’s cognitive profile. All of them are roots of the graph. A connection $A\rightarrow B$ denotes that $B$ is conditionally dependent on $A$. Linking functions are used to formally define relationships between dependent nodes. Nodes can encode meta-features, latent capabilities, and derived nodes such as observable performance or Intermediate Non-Observable Nodes (INON) for representing intermediate marginal performances.
All three node types were seen in Fig. 1(b). The connective structure can be enriched by domain knowledge determining which cognitive profile elements and meta-features are expected to affect the values of instance-level inference nodes. These relationships can be higher-order and complex, with certain INONs requiring complex combinations of many cognitive profile elements and meta-features. The need for complexity can arise from the many confounding behaviours that a well-designed benchmark needs to account for. By a linking function, we refer to any differentiable function that maps the values from the outputs of one (or multiple) node’s probability distribution to the input of another. For example, it could be a sigmoid function of the difference between capabilities and demands. The functions can scale the incoming nodes, so that we adjust the ranges for capabilities, biases and noise. We give the specific equations for the linking functions (and node formulations) that we used in this work in Tables 11 and 3 in the appendix.
@@
A key conceptual difference between our measurement layouts and typical uses for HBNs is that our approach is trying to capture a hierarchical dependency relation on capabilities and demands, rather than encoding information on hyper-priors. The linking functions between nodes are often non-linear, conceived from domain knowledge of the task and applying ideas from cognitive modelling. One of the key decisions is to determine how capabilities interact and the extent to which they are compensatory, being realised with additive or multiplicative expressions.
We use PyMC (Salvatier et al., 2016b) to implement and fit measurement layouts in practice using the No U-Turn Sampler (Hoffman and Gelman, 2014), although any probabilistic programming language can be used for this.
3 Experiment 1: A Simple Navigation Task
Figure 2: A: The average final score of each agent, varying pixel input size and navigational noise. B: A measurement layout relating two task demands and a capability to performance. C: The Brier Scores on the held-out test set of performances, as predicted by the average performance on the train set (Aggregate; Orange) and the measurement layout (Model; Blue), against the expected Brier Score on the test set (orange line). The measurement layout outperforms the aggregate for every agent. Note: No agent passed more than around 50% of the instances.
To demonstrate the power of the measurement layout approach, we design a simple experiment with agents with predefined capabilities. These agents interact with a three-dimensional embodied environment, Animal-AI (Beyret et al., 2019; Voudouris et al., 2025), and can move forwards, backwards, and turn; their goal is to obtain a green reward in each episode. They receive an $n\times n$ RGB image, and turn around until there is a green pixel directly ahead, and then they move towards it. We can vary the resolution of their image observations, directly manipulating their capability to see, i.e., their visual acuity. We can also vary the speed with which they move towards green rewards, by introducing random stationary actions once they have located a green reward, directly manipulating their navigation capability. We generate 1000 agents with different visual acuities and navigation capabilities, and test them on 1000 tasks in Animal-AI consisting of green rewards of different sizes placed at different distances from the agent. The average final reward across these 1000 tasks is presented in Figure 2A, demonstrating the effect of manipulating the size of the observation image and the amount of navigational noise.
Figure 2B presents the measurement layout for this task, relating the goal distance and the goal size to two latent capabilities: visual acuity and navigation capability. We relate visual acuity to the quantity $\frac{\text{distance}}{\text{size}}$. Smaller, more distant rewards are harder, and therefore require higher visual acuity, than larger, closer rewards. Navigation capability is only related to goal distance. We use half-normal priors on these latent capabilities, reflecting the fact that there is a meaningful 0 value for both navigation capability and visual acuity, representing no capability. We fitted the measurement layout in Figure 2A using 4 chains, with 1000 warm up draws and 2000 sample draws to ensure convergence. Figure 2C shows how our measurement layout outperforms the aggregate as a predictor of success for all agents. Figure 3 shows that the inferred posterior estimates of the capability neatly correspond to the relative visual acuity and navigation capability of a sample of agents.
Figure 3: Abilities for a selected subset of vision agents, with means (points) and 95% Highest Density Intervals. Left: The navigation abilities for agents with $\times32$ and $\times40$ pixel inputs, for noise levels of 0.0, 0.3, 0.8, and 0.9. Differences in navigation are recovered given a pixel input size. Right: The visual acuities for agents with $\times4$, $\times20$, and $\times40$ pixel inputs, for low (0.0) and high (0.8) navigational noise. Differences in visual acuity are recovered given a navigational noise level. Navigation ability is dominated by visual acuity because the agent can only navigate (with noise) to a goal that it can see.
4 Experiment 2: An Object Permanence Task
Our second experiment explores a more complex set of capabilities, including Object Permanence. To ensure that we knew the ground truth abilities of the agents, we created a synthetic dataset by simulating the performance of 30 synthetic agents based on specified cognitive profiles. To achieve this, the team was split into Team A who designed the synthetic agents, and Team B who built the measurement layout without knowledge of the agents built by Team A. More details of the specific testbed and the synthetic agents can be found in Appendix C.2. We present the measurement layout for the object permanence tasks in Fig. 4. In the object permanence benchmark (Voudouris et al., 2022), there are additional challenges to navigation in the form of ramps, platforms, and lava. Subsequently, meta-features for the presence of these challenges, as well as nodes for the agent’s capability (and instance-level performance) at handling them must be added to the measurement layout. In order to represent object permanence specifically, we add extra meta-features corresponding to key OP demands. A full mathematical description is given in Appendix D.2